OpenStudy (anonymous):

can someone explain to me how they get this : ∑_(r=1)^n▒u_r = ∑_(r=1)^n▒r(r+1) = r(r+1)(r+2) - (r-1)r(r+1) by using the method of difference , I stil don't get it why they end up like that,,,

OpenStudy (cruffo):

I'm not seeing the problem correctly, looks like some things are missing. could someone type this out .

OpenStudy (cruffo):

\[\large \sum_{r=1}^n ? u_r = \sum_{r=1}^n ?r(r+1) = (r+1)(r+2) - (r-1)r(r+1)\]

OpenStudy (anonymous):

i want to ask my own questions

OpenStudy (anonymous):

this problem same as this question : Find the sum of n terms of the series 1.4.7 + 4.7.10 + 7.10.13 \[U_{r}\] = (3r-2)(3r+1)(3r+4) Form another sequence Vr by adding one more factor to the end of the general term Ur \[V_{r}\] = (3r-2)(3r+1)(3r+4)(3r+7) \[V_{r-1}\] = (3r-5)(3r-2)(3r+1)(3r+4) my question is : how they get the another factor in the general term & add it to another sequence ? In this case for \[V_{r}\] they get the another factor is (3r+7) and for \[V_{r-1}\] they get (3r-5)...HOW...????